# Characterizing NGC 6383: A study of pre-main sequence stars, mass segregation, and age using Gaia DR3 and 2MASS

### v1

@@ 2.3 Sagitta
Sagitta is a neural network designed to classify stars as pre-main sequence (PMS) and estimate their ages, trained with data from Gaia DR2 and the 2MASS. It incorporates three convolutional neural networks each dedicated to determining stellar extinction ($A_{v}$), PMS probability, and estimating stellar ages. Sagitta’s age estimation is applicable for young stars up to $\sim 80~{}\mathrm{Myr}$ and relies on inputs including Gaia parallaxes, and average line-of-sight extinction (McBride et al., 2021).
This tool’s efficacy stems from its training on a well-curated dataset of PMS stars within moving groups (Kounkel et al., 2020), and is supplemented by various literature sources that have provided previously measured ages. Sagitta uniquely offers stability in age predictions across G, K, and M-type stars within the same population, surpassing traditional isochrone techniques in consistency (Kounkel et al., 2023).
By identifying PMS stars and their ages within NGC 6383, it allows for a nuanced understanding of the cluster’s star-forming history. It is important to note that Sagitta is most reliable for PMS stars; the photometry of higher-mass stars that have reached the main sequence is not indicative of their age, limiting the model’s applicability to stars that remain in the PMS stage.
Figure 9: Normalized cumulative counts of stars within NGC 6383 are plotted against radial distance from the cluster center for different brightness ranges, segmented into quartiles of $G_{\text{mag}}$ from the brightest at 8.80 mag to the faintest at 20.66 mag. The green dashed line at $R_{c}=1.94$ arcmin marks the core radius, and the red dashed line at $R_{t}=40.71$ arcmin denotes the tidal radius, illustrating the spatial distribution of stars with varying brightness within these boundaries. See the text for more reference.
@@ 3 Results
In this section, we report the primary findings of our study, focusing on the mode of the distributions for members with a membership probability of at least $60$ per cent.
Table 2 summarizes the derived parameters for sources with membership probabilities exceeding $60$ per cent. The probability distribution for these sources is illustrated in Fig. 1. The condensed cluster tree is depicted in Fig. 17, and Fig. 2 indicates the cluster’s center. The proper motions of the cluster members are shown in Fig. 3, while Fig. 4 shows the distance and parallax estimations. The structural parameters of the cluster, alongside the fitted King Profile, are presented in Fig. 5. Fig. 6 provides a visualization of the tidal radius, core radius, half-light radius, half-mass radius, the Hill radius, and the gravitational bound radius. Finally, Fig. 7 displays the best-fit isochrone and the Color-Magnitude Diagram (CMD) of NGC 6383.
Applying the HDBSCAN algorithm returned several clusters, predominantly composed of non-physical clusters. However, one significant cluster is recognized, distinguished by its proper motions. This cluster contained $701$ sources initially, and after applying a $2\sigma$ clipping based on the methodology outlined, $321$ sources were retained.
Figure 10: Histograms of parameters obtained with Sagitta for NGC 6383 potential members with a PMS probability over 60 percent. Upper panel: Distribution of PMS probability values, indicating the likelihood of stars being pre-main-sequence stars. Middle panel: Distribution of logarithmic age values, illustrating the range and frequency of star ages within the cluster. Lower panel: Distribution of visual extinction values ($A_{V}$), representing the amount of dust extinction affecting the observed stars.
@@
The central position of the cluster in equatorial coordinates is $263.6826\pm 0.1122\,^{\circ}$ in RA and $-32.5838\pm 0.1122\,^{\circ}$ in DEC. These errors are calculated from the quadratic sum of the mean errors in RA and DEC and the optimal KDE bandwidth. The mean proper motion values are $2.5419\pm 0.0094\,\mathrm{mas\,yr^{-1}}$ in RA and $-1.7127\pm 0.0086\,\mathrm{mas\,yr^{-1}}$ in DEC, with dispersions of $0.1532$ and $0.1383$ respectively, which align with the acceptable values from Perren et al. (2023). The projected velocity is calculated as $3.068\pm 0.009\,\mathrm{mas\,yr^{-1}}$.
@@
The mean parallax derived is $0.908\pm 0.004\,\mathrm{mas}$, with a mean sampled distance of $1.107\pm 0.035\,\mathrm{kpc}$. These values slightly deviate from the results obtained by earlier studies (Trumpler, 1930; Sanford, 1949; Fitzgerald et al., 1978; The et al., 1985; Paunzen et al., 2007) but are consistent with more recent findings (Becker & Fenkart, 1971; Kharchenko et al., 2005; Piskunov et al., 2008; Jaehnig et al., 2021; He et al., 2022; Pulgar-Escobar & Henríquez-Salgado, 2024; Hunt & Reffert, 2024).
@@
Radial velocities are estimated with a median of $-0.67\pm 33.056\,\mathrm{km}\,\mathrm{s}^{-1}$ using $16$ clustered sources with probabilities above $60$ per cent and a binary probability lower than $60$ per cent. The substantial standard deviation primarily stems from the limited availability of radial velocity data—only $6.299$ per cent of our Gaia-selected sources. Figure 15 illustrates the sources with available radial velocities, their magnitudes, and the amplitude of their radial velocities.
Figure 11: Upper Panel: Color-magnitude diagram displaying probable members and members of NGC 6383, color-coded by total mass, which ranges from lower to higher mass (red to blue). The dashed black line represents the best mean fit isochrone, and the dotted black line represents the best median fit isochrone, illustrating the evolutionary track inferred for the cluster. The total mass is calculated as the sum of the two components if the binary probability is greater than 0.7. Lower Panel: Same stars as in the upper panel, now color-coded by binary probability, illustrating the likelihood of binary systems within these stars from low to high (red to blue).
@@
Figure 12: Left panel: Cumulative distributions of single stars within NGC 6383 are segmented into four mass quartiles ranging from $0.23-0.69M_{\odot}$ to $1.52-13.08M_{\odot}$, illustrating the spatial distribution across different mass segments and highlighting the potential influences of mass on stellar distribution within the cluster. Middle panel: Similar distributions for binary stars are displayed. Right panel: The normalized cumulative distributions compare single versus binary stars, emphasizing differences within the cluster. All plots are based on stars with a membership probability of at least $60$ per cent. Kolmogorov-Smirnov test results are included to quantitatively assess the differences in distributions, underscoring the statistical validity of observed segregation patterns. For a detailed discussion of the Kolmogorov-Smirnov test results, please refer to the main text.
@@ 3.1 Structural Parameters
The cluster exhibits a tidal radius of $40.712\pm 14.379\,\mathrm{arcmin}$. The core radius, indicating the densest region, is $1.943\pm 0.1857\,\mathrm{arcmin}$. The parameters $k$ and $b$, representing the King model structural constants, are $4.902\pm 0.413$ and $0.011\pm 0.0062$ stars per square arcminute, respectively. These results are visualized in Fig. 5. The concentration parameter $C$, defined as $\log(R_{t}/R_{c})$, is approximately $3.042$, calculated following Peterson & King (1975). The half-light radius is determined to be $6.023\,\mathrm{arcmin}$.
The cluster’s outer limits extend to approximately $40.236\,\mathrm{arcmin}$, as shown in Fig. 2. We identified $38$ sources between the Hill radius and the tidal radius, with membership probabilities ranging from $0.6$ to $1$, averaging $0.82$. These sources, depicted in Fig.6, are likely not gravitationally bound to the cluster due to the extensive area covered by the cone search, implying they share similar proper motion but are not necessarily physically associated with the cluster.
@@ 3.2 Age, Extinction, and Color-Magnitude Diagram
The Color-Magnitude Diagram (CMD) of NGC 6383 is displayed in Fig. 7, including Pre-Main Sequence (PMS) stars and the median and mean isochrone. The optimal model suggests a logarithmic age of $6.54\pm 0.155$, a distance modulus of $10.416\pm 0.748$ corresponding to a distance of $1.211_{0.858}^{1.709}\,\mathrm{kpc}$, and a metallicity $Z$ of $0.023\pm 0.008$. An extinction correction of $A_{V}=1.221\pm 0.241$ was applied. The model’s modulus distance aligns with our parallax-derived estimates.
The CMD, particularly sources brighter than G magnitude of $12\,\mathrm{mag}$, displays a well-defined main sequence. The presence of PMS stars and lower-mass main sequence stars broadens the lower part of the CMD, complicating the fit of a single isochrone. Figure 8 presents various isochrones for Gaia and 2MASS data, considering ages between $6.2$ and $7.0$ in logarithmic scale and initial metallicities $Z$ of $0.023$. The age range of star formation in NGC 6383 appears to span from $6.2$ to $6.8$ in logarithmic scale.
@@ 3.3 Identification of YSOs
Our analysis aims to quantify the fraction of Young Stellar Objects (YSOs) within the cluster, noted as $Y_{\text{frac}}$, which serves as an age indicator. We determined the reddening-free parameter $Q$ for each star using the formula:
$Q=(J-H)-\frac{E(J-H)}{E(H-K)}(H-K)$ (4)
where $E(J-H)/E(H-K)$ is assumed to be $1.55$ (Mathis, 1990). Following Buckner & Froebrich (2013), a star is considered a YSO if its $Q$ value is less than $-0.05\,\mathrm{mag}$.
To estimate the lower limit of $Y_{\text{frac}}$, we assume the presence of at least $N_{\text{YSO}}-\sqrt{N_{\text{YSO}}}$ young stars, accounting for potential photometric scatter and the typically small numbers of YSOs. The fraction is calculated as
$Y_{\text{frac}}=\frac{N_{\text{YSO}}-\sqrt{N_{\text{YSO}}}}{N_{\text{cl}}},$ (5)
where $N_{\text{YSO}}$ is the number of YSOs $\sqrt{N_{\text{YSO}}}$ represents the uncertainty due to potential photometric scatter, and $N_{\text{cl}}$ is the total number of cluster members.
We identified $53$ YSOs, with a mean $Q$ value of $-0.52$, a median of $-0.38$, and a standard deviation of $0.42$ for sources below $-0.05\,\mathrm{mag}$. The resulting $Y_{\text{frac}}$ is $0.18$, which is higher compared to clusters analyzed in Buckner & Froebrich (2013).
Figure 14: Left panel: Same as Fig. 12, with cumulative distributions of single stars within NGC 6383. Middle panel: Similar distributions for binary stars. Right panel: The normalized cumulative distributions compare single versus binary stars. These plots were generated with data filtered for a minimum segment mass cutoff of $2.92M_{\odot}$, corresponding to a segregation time of 4 Myr. For a detailed discussion of the K-S test results, please refer to the main text.
@@ 3.4 Luminosity, Mass, and Dynamical State
The luminosity function (LF) represents the distribution of a cluster’s stars across different magnitude bins. For NGC 6383, we converted apparent $G$ magnitudes to absolute magnitudes using the parallax-derived distance. Figure 13 shows the LF, indicating an increase towards dimmer magnitudes, alongside the luminosity function derived from apparent magnitudes.
@@
The radial cumulative distributions of cluster members at various magnitude levels (Fig. 9) show that fainter stars ($17.93-20.66\,\mathrm{mag}$) are less concentrated around the cluster’s center. This observation is consistent with Gaia’s completeness limit for magnitudes between $18$ and $19$ at distances around $1\,\mathrm{kpc}$ (Boubert & Everall, 2020; Castro-Ginard et al., 2023). The significance of this distribution was confirmed using a Kolmogorov-Smirnov (K-S) test, yielding p-values from $0.009$ to $0.052$ when compared with other ranges.
@@
AsTeCA calculates the probability of a star being part of a binary system and estimates the masses of the primary and secondary stars. However, these estimates are less reliable in the PMS stage due to the poorly defined CMD, as previously discussed. Figure 11 displays these binary probabilities and total mass, indicating that these values should be interpreted with caution, as they may not be entirely accurate. The average mass of the stars is $1.26\,M_{\odot}$.
@@
To investigate mass segregation, we constructed radial normalized cumulative mass distributions for single stars, binary stars, and both combined (Fig. 12). The K-S test was employed to compare these distributions, indicating a significant difference between binary and single stars, with a p-value of $0.026$ and a K-S test statistic of $0.2079$. This suggests binary mass segregation, where binary stars are more centrally concentrated. Additionally, the bin mass range between $1.52-13.08~{}M_{\odot}$ shows a significant difference compared to the other mass ranges, with a p-values ranging from $0.0683$ to $0.0840$ and a K-S test statistic of $\sim 0.2760$, but this mass range is broad and cannot ensure mass segregation within it.
@@
To contextualize these findings, we calculated the half-mass relaxation time ($t_{\mathrm{rh}}$):
$t_{\mathrm{rh}}=\frac{0.17N}{\ln(\lambda N)}\sqrt{\frac{r_{h}^{3}}{GM}}$ (6)
@@
where $N$ is the number of cluster members, $\lambda$ is a constant according to Giersz & Heggie (1994), and $r_{\mathrm{hm}}=1.65\,\mathrm{pc}$ is the half-mass radius, derived from the radial cumulative mass distribution. The calculated half-mass relaxation time for NGC 6383 is $9.808\,\mathrm{Myr}$.
@@
For estimating the minimum mass segregation time ($t_{\mathrm{seg}}$) for the cluster’s most massive star, we follow the approach of Spitzer (1969):
$t_{\mathrm{seg}}=\frac{\langle m\rangle}{m}t_{rh}$ (7)
@@
where $\langle m\rangle$ is the average mass of a star in the cluster, and $m$ is the mass of the most massive star. This results in a segregation time of $1.364\,\mathrm{Myr}$ for the most massive star of $13.08\,M_{\odot}$, indicating it has had sufficient time to undergo mass segregation.
@@
Finally, we evaluated the potential for primordial mass segregation by comparing the radial distributions of stars with $t_{seg}>4\,\mathrm{Myr}$ to the cluster’s age of approximately $4\,\mathrm{Myr}$. The sample includes stars with masses ranging from $0.22\,M_{\odot}$ to $3.42\,M_{\odot}$, and the analysis revealed a similar behavior as the Fig. 12, the highest mass bin has a different distribution. In this case, the mass range is not broad, with p-values of $0.0653$ to $0.3364$ and a K-S test statistic of $0.1976$ to $0.2697$, which indicates primordial mass segregation of the higher mass single sources between the range of $1.39\,M_{\odot}$ and $3.42\,M_{\odot}$. In this sample, the average mass is $1.05\,M_{\odot}$. The distributions of single and binary stars are similar up to approximately $5\,\mathrm{arcmin}$, beyond which single stars show less concentration compared to binary stars.

### v2

@@ 2.3 Sagitta
Sagitta is a neural network designed to classify stars as PMS and estimate their ages. It is trained with data from Gaia DR2 and the 2MASS. It incorporates three convolutional neural networks each dedicated to determining stellar extinction ($A_{v}$), PMS probability, and estimating stellar ages. Sagitta’s age estimation is applicable for young stars up to $\sim 80\leavevmode\nobreak\ \mathrm{Myr}$ and relies on inputs including Gaia parallaxes, and average line-of-sight extinction (McBride et al., 2021).
The tool’s efficacy stems from its training on a well-curated dataset of PMS stars within moving groups (Kounkel et al., 2020), which is supplemented by various literature sources with previously measured ages. Sagitta uniquely offers stability in age predictions across G, K, and M-type stars within the same population, surpassing traditional isochrone techniques in consistency (Kounkel et al., 2023).
The identification of PMS stars and the estimation of their ages within NGC 6383, allows for a detailed understanding of the cluster’s star formation history to be gathered. It is important to note that Sagitta is most reliable for PMS stars, as the photometry of higher-mass stars that have reached the main sequence is not indicative of their age, limiting the model’s applicability to stars that remain in the PMS stage.
@@ 3 Results
In this section, we report the results of our study, focusing on the mode of the distributions for members with a membership probability of at least $60$ percent.
Table 2 summarizes the parameters derived for sources with membership probabilities exceeding $60$ percent. The probability distribution for these sources is illustrated in Fig. 1. The condensed cluster tree is depicted in Fig. 17, and Fig. 2 indicates the cluster’s center. The proper motions of the cluster members are shown in Fig. 3, while Fig. 4 shows the distance and parallax estimations. The structural parameters of the cluster, alongside the fitted King Profile, are presented in Fig. 5. Fig. 6 provides a visualization of the tidal radius, core radius, half-light radius, half-mass radius, the Hill radius, and the gravitational bound radius. Finally, Fig. 7 displays the best-fit isochrone and the CMD of NGC 6383.
Applying the HDBSCAN algorithm returned several clusters. A major flaw of HDBSCAN, as shown in Hunt & Reffert (2021), is its high false positive rate. This is due to the algorithm’s overconfidence, often reporting dense random fluctuations of a given dataset as clusters. For the problem of the determination of the membership, the proper motion of the stars was used to assess the significance of an identified cluster. As expected for the case of NGC6383, one significant cluster was identified. This cluster initially contained 701 sources, and after applying the $2\sigma$ clipping described in Section 2.1.1, 321 sources were retained.
Figure 11: Upper Panel: CMD displaying probable members and members of NGC 6383, color-coded by total mass, ranging from lower (red) to higher mass (blue). The dashed black line represents the best mode fit isochrone, illustrating the evolutionary track inferred for the cluster. The total mass is calculated as the sum of the two components if the binary probability is greater than $0.7$. Black stars represent members and probable members with no available 2MASS data, and thus no calculated mass. Lower Panel: The same stars as in the upper panel, now color-coded according to binary probability, illustrating the likelihood of binary systems within these stars, from low (red) to high (blue). The black dashed line again represents the best mode fit isochrone. Black stars represent members and probable members with no 2MASS data, thus lacking binary probability estimates.
Figure 13: Normalized cumulative counts of stars within NGC 6383 are plotted against radial distance from the cluster center for different brightness ranges, segmented into quartiles of $G_{\text{mag}}$ from the brightest at $8.80$ mag to the faintest at $\sim 20.7$ mag. The green dashed line at $R_{c}=1.95\,\mathrm{arcmin}$ marks the core radius, and the red dashed line at $R_{t}=40.4\,\mathrm{arcmin}$ denotes the tidal radius, illustrating the spatial distribution of stars with varying brightness within these boundaries. See the text for more reference.
@@
The central position of the cluster in equatorial coordinates is $263.6826\pm 0.1122\,^{\circ}$ in RA and $-32.5838\pm 0.1122\,^{\circ}$ in DEC. These errors are calculated from the quadratic sum of the mean errors in RA and DEC and the optimal KDE bandwidth. The mean proper motion values are $2.5400\pm 0.0096\,\mathrm{mas\,yr^{-1}}$ in RA and $-1.7100\pm 0.0086\,\mathrm{mas\,yr^{-1}}$ in DEC, with dispersions of $0.153$ and $0.138$ respectively, which align with the acceptable values from Perren et al. (2023). The calculated projected velocity is $3.070\pm 0.010\,\mathrm{mas\,yr^{-1}}$.
@@
The mean parallax derived is $0.908\pm 0.004\,\mathrm{mas}$, with a mode sampled distance of $1.1100\pm 0.0600\,\mathrm{kpc}$. These values slightly deviate from the results obtained in earlier studies (Trumpler, 1930; Sanford, 1949; Fitzgerald et al., 1978; The et al., 1985; Paunzen et al., 2007) but are consistent with more recent studies (Becker & Fenkart, 1971; Kharchenko et al., 2005; Piskunov et al., 2008; Jaehnig et al., 2021; He et al., 2022; Pulgar-Escobar & Henríquez-Salgado, 2024; Hunt & Reffert, 2024).
@@
Radial velocities are estimated with a median, mean, and standard deviation of $-6.11$, $-15.1$, and $31.1\,\mathrm{km}\,\mathrm{s}^{-1}$, respectively, using 16 clustered sources with probabilities above $60$ percent and a binary probability lower than $60$ percent. The substantial standard deviation primarily stems from the limited availability of radial velocity data, only $11.4$ percent of our Gaia-selected sources. The radial velocity calculation can be affected by sources with a relatively high radial velocity amplitude, which could indicate unresolved binaries and introduce significant variability in the measurements. To address this, as mentioned before, we only calculated the radial velocity of the cluster using sources with a binary probability lower than $60$ percent. This ensures that the derived cluster velocity is more representative of the overall motion of the cluster’s stars, reducing the impact of outliers caused by binary systems. However, as mentioned, this value is still not fully representative of the cluster since the sample does not sufficiently capture the full dynamics of the cluster. Figure 9 illustrates all the probable members and members with available radial velocities, their magnitudes, and the amplitude of their radial velocities.
@@
Figure 14: Left panel: The cumulative distributions of single stars within NGC 6383 are segmented into four mass quartiles ranging from $0.360-0.720\,M_{\odot}$ to $1.82-14.7\,M_{\odot}$, illustrating the spatial distribution across different mass segments and highlighting the potential influences of mass on the distribution of stars within the cluster. Middle panel: The same distributions as in the left panel for binary stars only, where binaries are defined as sources with a binary probability greater than $60$ percent. Right panel: The normalized cumulative distributions compare single stars to binary stars, with masses between $0.360\,M_{\odot}$ and $14.7\,M_{\odot}$. All plots are based on stars with a membership probability of at least $60$ percent. Kolmogorov-Smirnov (KS) test results are detailed in Sec. 3.4, to quantitatively assess the differences between distributions, underscoring the statistical validity of observed segregation patterns.
@@ 3.1 Structural Parameters
The cluster exhibits a tidal radius of $40.4\pm 14.3\,\mathrm{arcmin}$ ($13.10\pm 4.71\,\mathrm{pc}$) and a core radius of $1.950\pm 0.190\,\mathrm{arcmin}$ ($0.63200\pm 0.00621\,\mathrm{pc}$). The King model parameters $k$ and $b$, representing structural constants, are $4.910\pm 0.437$ and $0.01110\pm 0.00630$ stars per square arcminute, respectively. The concentration parameter $C$, defined as $\log(R_{t}/R_{c})$, is $3.03$, following Peterson & King (1975). The half-light radius is determined to be $6.02000\pm 0.00060\,\mathrm{arcmin}$ ($1.96000\pm 0.00020\,\mathrm{pc}$) and the half-mass radius is $6.240\pm 0.251\,\mathrm{arcmin}$ $(1.650\pm 0.066\,\mathrm{pc})$. Part of these results are represented in Fig. 5.
The cluster’s outer limits extend to approximately $40.2\,\mathrm{arcmin}$ ($13.1\,\mathrm{pc}$), as shown in Fig. 2. We identified $39$ sources between the Hill radius and the tidal radius, with membership probabilities ranging from $0.6$ to $1.0$, averaging $0.82$. These sources, plotted in Fig.6, are likely not gravitationally bound to the cluster. This implies that they have proper motions similar to the cluster stars but are not necessarily physically associated with the cluster, due to the extensive area covered by the cone search.
@@ 3.2 Age, Extinction, and Color-Magnitude Diagram
The CMD of NGC 6383 is displayed in Fig. 7, including PMS stars and the median, mean and mode isochrones. The mode indicates a logarithmic age in years of $6.550\pm 0.145$, a distance modulus of $10.300\pm 0.262\,\mathrm{mag}$ corresponding to a distance of $1.150_{-0.130}^{+0.147}\,\mathrm{kpc}$, and a metallicity $Z$ of $0.024\pm 0.008$. An estimated extinction correction of $A_{V}=1.240\pm 0.262$ was applied. The model’s distance modulus aligns with our parallax-derived estimates.
The CMD, particularly for sources brighter than $G=12.0\,\mathrm{mag}$, displays a well-defined main sequence. The presence of PMS stars and lower-mass main sequence stars broadens the lower part of the CMD, complicating the fit of a single isochrone. Figure 8 presents various isochrones for Gaia and 2MASS data, considering ages between $6.20$ and $7.00$ in logarithmic scale and initial metallicities $Z$ of $0.024$. The age range of star formation in NGC 6383 appears to span from $6.20$ to $6.80$ in logarithmic scale.
@@
Figure 15: Left panel: Same as Fig. 14, with cumulative distributions of single stars within NGC 6383. Middle panel: The same distributions for binary stars. Right panel: The normalized cumulative distributions compare single versus binary stars. These plots were generated with data filtered for a minimum segment mass cutoff of $6.08\leavevmode\nobreak\ M_{\odot}$, corresponding to a segregation time of $3.53\leavevmode\nobreak\ \mathrm{Myr}$. For a detailed discussion of the K-S test results, please refer to the main text.
@@ 3.3 Identification of YSOs
Young stellar Objects (YSOs) are early-stage stars still accreting material from their surrounding disks. They are characterized by significant variability due to the accretion processes. YSOs include both protostars and PMS stars. PMS stars, which encompass both classical T Tauri stars and Herbig Ae/Be stars, represent a later stage in stellar evolution. Thus, while PMS stars are a type of YSO, they are more evolved than protostars.
Our analysis aims to quantify the fraction of YSOs within the cluster, noted as $Y_{\text{frac}}$, which serves as an age indicator. We determined the reddening-free parameter $Q$ for each star using
$Q=(J-H)-\frac{E(J-H)}{E(H-K)}(H-K),$ (4)
where $E(J-H)/E(H-K)$ is assumed to be $1.55$ (Mathis, 1990). Following Buckner & Froebrich (2013), a star is considered a YSO if its $Q$ value is less than $-0.050\,\mathrm{mag}$.
To estimate $Y_{\text{frac}}$ more rigorously, we employed a Bayesian approach for determining confidence intervals for binomial proportions, which provides a more statistically robust estimate compared to traditional methods (Cameron, 2011). Using a non-informative $\mathcal{B}(1,1)$ prior, the posterior distribution for $Y_{\text{frac}}$ is modeled as
$Y_{\text{frac}}=\frac{N_{\text{YSO}}}{N_{\text{cl}}},$ (5)
where $N_{\text{YSO}}$ is the number of YSOs and $N_{\text{cl}}$ is the total number of cluster members. The $95\%$ credible interval for $Y_{\text{frac}}$ was then derived from the posterior Beta distribution.
We identified 53 YSOs, with a mean $Q$ value of $-0.520$, a median of $-0.380$, and a standard deviation of $0.420$ for sources below $-0.050\,\mathrm{mag}$.
The resulting $Y_{\text{frac}}$ is $0.280$, with a $95\%$ credible interval of [$0.220$, $0.340$]. This value is higher compared to clusters analyzed in Buckner & Froebrich (2013), where only 18 of 397 cluster candidates in their sample had $Y_{\text{frac}}>0.100$, with no objects having $Y_{\text{frac}}>0.200$. In contrast, our results indicates a higher fraction of YSOs compared to their sample. This difference suggests that NGC 6383 is relatively younger or has a higher recent star formation activity compared to most clusters in their study.
@@ 3.4 Luminosity, Mass, and Dynamical State
The luminosity function (LF) represents the distribution of the magnitude of the cluster members. For NGC 6383, we converted apparent $G$ magnitudes to absolute magnitudes using the parallax-derived distance. Figure 12 shows the LF, indicating an increase towards dimmer magnitudes, alongside the luminosity function derived from apparent magnitudes.
@@
The radial cumulative distributions of cluster members at various magnitude levels, illustrated in Fig. 13, show that fainter stars, ranging from $17.93$ to $20.66\,\mathrm{mag}$, are less concentrated around the cluster’s center. This observation is consistent with Gaia’s completeness limit for magnitudes between $18$ and $19$ at distances around $1\,\mathrm{kpc}$ (Boubert & Everall, 2020; Castro-Ginard et al., 2023). The significance of this distribution was confirmed using a Kolmogorov-Smirnov (K-S) test, yielding p-values from $0.009$ to $0.052$ when compared with other ranges.
@@
ASteCA calculates the probability of a star being part of a binary system and estimates the masses of the primary and secondary stars. However, these estimates are less reliable in the PMS stage due to the poorly defined CMD, as previously discussed. Figure 11 displays these binary probabilities and total mass, indicating that these values should be interpreted with caution, as they may not be entirely accurate. The average mass of the stars is $1.59\,M_{\odot}$.
@@
To investigate mass segregation, we constructed radial normalized cumulative mass distributions for single stars, binary stars, and the combination of the two types of stellar systems, as shown in Fig. 14. The K-S test was employed to compare these distributions, indicating a significant difference between binary and single stars, with a p-value of $0.07$ and a K-S test statistic of $0.2079$. This suggests binary mass segregation, where binary stars are more centrally concentrated. Additionally, the last quartile mass range between $1.82$ and $14.71\leavevmode\nobreak\ M_{\odot}$ shows a significant difference compared to the other mass ranges, with p-values ranging from $0.04$ to $0.95$ and a K-S test statistic ranging from $0.14$ to $0.37$. However, this mass range is broad and cannot ensure mass segregation within it.
@@
To contextualize these findings, we calculated the half-mass relaxation time $t_{\mathrm{rh}}$ (Spitzer, 1969), which is crucial for estimating the minimum mass segregation time $t_{\mathrm{seg}}$. The half-mass relaxation time is given by:
$t_{\mathrm{rh}}=\frac{0.17N}{\ln(\lambda N)}\sqrt{\frac{r_{hm}^{3}}{GM}}$ (6)
@@
where $N$ is the number of cluster members, $\lambda=0.110$ is a constant (Giersz & Heggie, 1994), and $r_{\mathrm{hm}}=1.650\pm 0.066\,\mathrm{pc}$ is the half-mass radius, derived from the radial cumulative mass distribution. The calculated half-mass relaxation time for NGC 6383 is $13.60\pm 3.87\,\mathrm{Myr}$.
@@
Using the half-mass relaxation time, we can estimate $t_{\mathrm{seg}}$ for the cluster’s most massive star by following the approach of Spitzer (1969):
$t_{\mathrm{seg}}=\frac{\langle m\rangle}{m}t_{\mathrm{rh}}$ (7)
@@
where $\langle m\rangle$ is the average mass of a star in the cluster, and $m$ is the mass of the most massive star. This results in a segregation time of $1.470\pm 0.418\,\mathrm{Myr}$ for the most massive star of $14.7\,M_{\odot}$, indicating it has had sufficient time to undergo mass segregation.
@@
We evaluated the potential for primordial mass segregation by comparing the radial distributions of stars with a $t_{\mathrm{seg}}$ greater than $3.53\,\mathrm{Myr}$, which corresponds to the age of the cluster. The sample includes stars with masses ranging from $0.360\,M_{\odot}$ to $6.08\,M_{\odot}$, which corresponds to $97.0$ percent of the original sample. The results, shown in Fig. 15, revealed a similar behavior to Fig. 14, where the highest mass bin has a distinct distribution. This similarity is expected, given that the sample is nearly identical to the one used in the previous analysis. The mass range is narrower compared to the previous figure, with p-values ranging from $0.0600$ to $0.450$ and a K-S test statistic from $0.210$ to $0.460$. This suggests that the higher mass single stars, between $1.39\,M_{\odot}$ and $3.42\,M_{\odot}$, do not exhibit the same cumulative growth as lower mass stars. In this sample, the average mass is $1.36\,M_{\odot}$. The distributions of single and binary stars are similar up to approximately $5\,\mathrm{arcmin}$, beyond which single stars show less concentration compared to binary stars.
Given the current dynamical state of the cluster and the number of stars (presumably larger at birth), it is unlikely that there has been sufficient time for dynamical mass segregation. If the cluster initially formed in a roughly spherical distribution without significant substructure, this suggests that binaries originated closer to the cluster center than single stars. Alternatively, simulations by McMillan et al. (2007) indicate that this outcome could result from star formation in smaller clumps that either formed mass-segregated or had short relaxation times, allowing for quick mass segregation before merging into the current cluster. These simulations conclude that clumps with initial mass segregation largely preserve this characteristic during merging. For clumps without initial mass segregation, dynamic mass segregation can occur before merging if their initial relaxation times are short, and this segregation is inherited by the merged cluster. Additionally, initial mass segregation and gas expulsion play a critical role in star cluster evolution (Marks et al., 2008). Regardless of the specific mechanism, this suggests a primordial, or near-primordial, central concentration of binaries during the formation process of the cluster, which is congruent with other young open clusters (Chen et al., 2007; Dib et al., 2007; Sabbi et al., 2008).